OpenStudy (deadshot):

What is the interval that represents the domain of the function f(x)=sin^(-1)x?

OpenStudy (deadshot):

\[f(x)=\sin ^{-1}x\]

ganeshie8 (ganeshie8):

whats the range of \(\sin(x)\) ?

OpenStudy (deadshot):

it is \[y=-\frac{ \pi }{ 2 }\le y \le \frac{ \pi }{ 2 }\]

ganeshie8 (ganeshie8):

no, thats the range of \(\sin^{-1}x\)

OpenStudy (deadshot):

oh

ganeshie8 (ganeshie8):

range of \(\sin(x)\) is \(-1 \le x \le 1\)

ganeshie8 (ganeshie8):

which will be same as domain of \(\sin^{-1}x\)

OpenStudy (deadshot):

Oh, okay! Thanks!

ganeshie8 (ganeshie8):

u wlc :)